OpenStudy (anonymous):

Find derivative of ln[sec x] Ok... I know the derivative of ln x = 1/x ... But when it has a function inside of it, what do I do? do I still do 1/sec x then use the chain rule?

OpenStudy (johnweldon1993):

Yeah that looks right... So we would have \[\large \frac{1}{\sec(x)} \times \sec(x)\tan(x)\] right?

OpenStudy (anonymous):

Yeah... Still gotta remember those damn trig derivatives and the antiderivatives, and then derive them for the AP test! lol...

OpenStudy (anonymous):

Thanks for the answer, anything else before I give out the best response? like any exceptions or whatever?

OpenStudy (johnweldon1993):

Nope that's about it...but what do you get when you simplify that?

OpenStudy (kc_kennylau):

You only have to memorize \(\dfrac d{dx}\sin x=\cos x\) and \(\dfrac d{dx}\cos x=-\sin x\) and you can derive all of them using quotient rule.

OpenStudy (anonymous):

wait... just realized that isn't one of the answers... so maybe if I convert it in terms of cos and sin... since sec x is just 1/cos x, it would be \[(\cos x ) + \frac{ 1 }{ \cos x} * \tan x\] so just tan x, which is sinx / cox

OpenStudy (johnweldon1993):

Right...the simplification would be tan(x)

OpenStudy (johnweldon1993):

\[\large \frac{1}{sec(x)} \times \sec(x)\tan(x) = \frac{\sec(x)\tan(x)}{\sec(x)} = \tan(x)\]

OpenStudy (kc_kennylau):

It's not plus; it's multiply @Psychedelic

OpenStudy (anonymous):

got that, thanks!